OpenStudy (anonymous):

Solve: x2 − 2x − 24 = 0

OpenStudy (anonymous):

do u have to factor

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

\[x^2 - 2x - 24 = 0\]Find two factors of -24 that add up to -2

OpenStudy (anonymous):

-6 , 4

OpenStudy (anonymous):

x² - 2x - 24 = 0 x² - x = 24 + (- 1)² x² - x = 24 + 1 (x - 1)² = 25 x - 1 = ± 5 = x - 1 - 5, = x - 6 = x - 1 + 5, = x + 4 Answer: (x - 6)(x + 4)

OpenStudy (anonymous):

x = 4, x = 6 x = 4, x = −6 x = −4, x = 6 x = −4, x = −6

OpenStudy (anonymous):

those are the answer choices

OpenStudy (anonymous):

Alright. Then you can place them by splitting the x^2 term. \[x^2 - 2x - 24 = 0 \implies (x - 6)(x + 4) = 0\]Now solve for x by setting each equal to 0. x - 6 = 0 x + 4 = 0 Can you solve these two equations?

OpenStudy (anonymous):

6, -4?

OpenStudy (anonymous):

Yup. x = 6, -4 :)

OpenStudy (anonymous):

thank you both(:

OpenStudy (anonymous):

Glad to help :)

OpenStudy (anonymous):

yup np